[ 2/4 0-1;01/26; 1/204] 13) (20 pts) Find the inverse of the matrix from question 12.
Answer #1

[2/4 ; 1/ 6 ; 1/2 04 0 0 Let A= 1 6 6 O 2 4 0 4 1.5 - 1 A1=10 ग 6 0 4 5.5 (2)-1(-25) = ).25 . 5 0 O AB (A)= At lies .5 T fr5